# Fisher-Hartwig asymptotics for non-Hermitian random matrices

### v1

@@ 1 Introduction
The main Theorem 1.2 below generalizes the above moment formula to the exponential generating function of the electric potential, coupled with mesoscopic linear statistics, for general confining external potentials. A remarkable aspect of this general formula is the relevance of a log-jump probability measure on the boundary of the droplet (see Definition 1.1), with the following consequence: the Gibbs measure associated to the electric potential converges to a Gaussian multiplicative chaos measure, which is universal up to a global Gaussian shift inherited from the log-jump measure (see Theorem 1.8).
@@
where the Barnes function $\rm G$} is defined in Subsection 1.5. Applications of general Fisher-Hartwig asymptotics for Toeplitz determinants (including the case of jump singularities) are multiple, we refer for example to [DeiItsKraII] for the relevance to phase transitions in the Ising model. The main contribution of our paper is the analogue of (1.2) – a formula due to Harold Widom in 1973 – for free fermions in 2d, as explained below.
@@
where ${\rm d}M$ is the measure induced by the Lebesgue measure on all $N\times N$ complex matrices and $V:\mathbb{C}\to\mathbb{R}\cup\{+\infty\}$ be a confining potential. This induces the following probability measure on the eigenvalues of $M$, which also corresponds to the distribution of free fermions on $\mathbb{C}$:
@@
Before stating it, we introduce the log-jump measure. For a function $g$ with domain including $\bar{S}$, let $g^{S}$ denote the unique bounded harmonic extension of $g$ to $S^{\rm c}$, with $g^{S}=g$ on $\bar{S}$. Note that for this harmonic extension, $\lim_{|z|\to\infty}g^{S}(z)=g^{S}_{\infty}$ exists and
$|g^{S}(z)-g^{S}_{\infty}|=\operatorname{O}\big(|z|^{-1}\big),\quad|\nabla^{n}g^{S}(z)|=\operatorname{O}\big(|z|^{-n-1}\big)\quad\textnormal{for all}\ n\geqslant 1\ \textnormal{as}\ |z|\to\infty.$
@@
Definition 1.1.
Let $\zeta$ be an arbitrary point in $S$ and $\log^{\zeta}=\log|\zeta-\cdot\,|$. We define the log-jump measure $\sigma=\sigma_{S}$ on the boundary of the droplet (${\rm d}s$ is the length measure on $\partial S$) as
$\sigma(z){\rm d}s=\frac{1}{2\pi}\mathcal{N}(\log^{\zeta})(z){\rm d}s,\quad z\in\partial S.$ (1.4)
It is a probability measure independent of the choice of $\zeta\in S$ (see Appendix B). We then define
$\mathfrak{s}=\int_{\partial S}\log^{\zeta}(z)\sigma(z){\rm d}s$
and this quantity is likewise independent of the choice of $\zeta$ (see Appendix B).
We also need the notations $L=\log\frac{\Delta V}{4}$, $L_{\infty}^{S}=\lim_{|z|\to\infty}L^{S}(z)$, and refer to Subsection 1.5 for the definition of the functional space $\mathscr{S}_{n,C,\kappa}$, to state the following main theorem. It gives the asymptotics for the joint Laplace transform of the electrostatic potential of free fermions, or equivalently the joint moments of the characteristic polynomial of random normal matrices.
@@
Let $C>1$, $m\in\mathbb{N}$, $0<\kappa\leqslant\frac{1}{100C}$ be fixed constants. Then, uniformly in $f\in\mathscr{S}_{5,C,\kappa}$, $\zeta_{1},\dots,\zeta_{m}\in\{\zeta\in S:\operatorname{dist}(\zeta,\partial S)>N^{-1/2+\kappa}\}$ satisfying $\min_{1\leqslant j<k\leqslant m}|\zeta_{j}-\zeta_{k}|>N^{-1/2+\kappa}$ and $\gamma_{1},\dots,\gamma_{m}\in[0,C]$ we have
$\mathbb{E}\Big[e^{\sum_{i=1}^{N}f(z_{i})}\prod_{j=1}^{m}\prod_{i=1}^{N}|z_{i}-\zeta_{j}|^{\gamma_{j}}\Big]=e^{N\int f{\rm d}\mu_{V}+\frac{1}{8\pi}\int_{\mathbb{C}}(|\nabla f^{S}|^{2}+\mathds{1}_{S}\Delta f+\Delta fL^{S}){\rm d}m}\prod_{j=1}^{m}e^{\frac{\gamma_{j}}{2}\big(\int_{\partial S}f\sigma{\rm d}s-f(\zeta_{j})\big)}$
$\times\prod_{j=1}^{m}e^{\gamma_{j}N\int\log^{\zeta_{j}}{\rm d}\mu_{V}}N^{\frac{\gamma_{j}^{2}}{8}}\frac{(2\pi)^{\frac{\gamma_{j}}{4}}}{{\rm G}\left(1+\frac{\gamma_{j}}{2}\right)}e^{\frac{\gamma_{j}}{4}(L(\zeta_{j})-L_{\infty}^{S})+\frac{\gamma_{j}^{2}}{8}(L(\zeta_{j})+2\mathfrak{s})}$
$\times\prod_{1\leqslant j<k\leqslant m}|\zeta_{j}-\zeta_{k}|^{-\frac{\gamma_{j}\gamma_{k}}{2}}e^{\frac{\gamma_{j}\gamma_{k}}{2}\mathfrak{s}}\left(1+\operatorname{O}(N^{-\kappa/10})\right)$
where $\rm G$} stands for the Barnes $\rm G$}-function (see Subsection 1.5).
The Ginibre ensemble is the paradigmatic non-Hermitian random matrix model; see [ByuFor] for its many facets. For this ensemble, our asymptotics take a simpler form. We have $\int_{\mathbb{D}}\log^{\zeta}\frac{{\rm d}m}{\pi}=\frac{|\zeta|^{2}-1}{2}$ and choosing $\zeta=0$ gives $\mathcal{N}(\log^{\zeta})=1$ and $\int_{\partial\mathbb{D}}\log^{\zeta}{\rm d}s=0$, so $\mathfrak{s}=0$. Hence the main theorem takes the following simplified expression, conjectured by Webb and Wong in [WebWong2019].
@@
To conclude this subsection, we note that moments of random characteristic polynomials of wide classes of random matrices have been a topic of major interest, see e.g. [AkeVer, FyoSup, BorStr, ForRai2009, BaiKea, Afa2019, SerSim2025] in the case of integer exponents by algebraic and supersymmetric methods, and [Kra2007, ItsKra, Cha, BerWebWon2017, WebWong2019, ChaFahWebWon, CFLW2021, ByuSeoYan, DeaMclMolSim2025] for fractional exponents by Riemann-Hilbert methods. Motivations for the study of random determinants in the non-Hermitian setting include for example the stability of large complex systems [BenFyoKho2021]. Theorems 1.2 and 1.3 introduce joint fractional moments for non-Hermitian ensembles, motivated by two-dimensional random geometry from free fermions, as explained below.
@@
In the 2d ambient space the Liouville quantum gravity measure also appears, for systems of free fermions, as stated in Theorem 1.8 below. The key ingredient is the precise asymptotics of the moments of the characteristic polynomial, Theorem 1.2 above. To state our result, note that from Corollary 1.7, the log-characteristic polynomial of a random normal matrix is $\frac{1}{2}$-log-correlated, so the corresponding critical value becomes $\gamma_{\rm c}=2\sqrt{2}$. The regime $\gamma<2$ is often referred to as the $L^{2}$-phase, while the full subcritical range $\gamma<2\sqrt{2}$ is called the $L^{1}$-phase. All arguments below are carried out in the $L^{1}$-phase.
@@
$\mathbb{E}\big[\log|\det(M-z)|\big]=N\int\log^{z}{\rm d}\mu_{V}+\frac{1}{4}+\frac{1}{4}(L(z)-L^{S}_{\infty})+\operatorname{O}(N^{-\varepsilon}).$
@@
where the second equality is shown in Appendix B. Let $\nu_{\gamma,(\mathfrak{s})}$ be the GMC measure associated to $X$, formally given by
@@
Our main theorems also allow to recover the leading order asymptotics for the electric field for free fermions and general potential: For $K$ any compact subset of $S$ with non-empty interior and any $\varepsilon>0$,
@@
Steps (2) and (4) rely on the method of Ward identities, or loop equations, which is common in field theories. For Hermitian random matrices, the seminal work [Joh1998] relied on loop equations to reduce Laplace transform estimates to a shift of the density of states. In our non-Hermitian setting, Ward identities are also instrumental as in [AmeHedMak2015], but the control of the error due to potential anisotropy (i.e. local directional contribution from the interactions) and the boundary of the droplet, is delicate for measures biased by determinant powers, i.e. in the presence of root singularities. Regarding the proof of local isotropy – a problem naturally absent in the settings of [BouFal2025, Keles2025] – Lemma 2.14 first treats nonsingular potentials, by a moments method specific to determinantal point processes but flexible enough to cover mesoscopic observables up to the boundary of the droplet; this estimate is then transferred to singular setting, as it holds with overwhelming probability (here $1-e^{-(\log N)^{D}}$ for any $D>0$), while in [AmeHedMak2015] angular cancellations hold in expectation for the nonsingular potentials. About the effect of the boundary, one novelty of our analysis is a careful control of the norms of the harmonic extensions of functions on mesoscopic scales, and an analysis of the effect of bulk singularities on the boundary field. This analysis is significant due to the log-jump measure, which generates an additional Gaussian shift when compared to the free field with free boundary conditions, harmonically extended outside of the droplet.
@@ Acknowledgements.
P. B. and A. K. were supported by the NSF standard grant DMS-2348202. G. D. gratefully acknowledges support from the Fondation de l’École Polytechnique, as well as from the Agence Nationale de la Recherche (ANR-25-CE40-5672 and ANR-25-CE40-1380). The research of L. H. was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) in the framework to the collaborative research center ”Multiscale Simulation Methods for Soft-Matter Systems” (TRR 146) under Project No. 233630050 and Project-No. 443891315 within SPP 2265.
@@ 2 Preliminaries on nonsingular potentials
For any given function $f$ of class $\mathscr{C}^{5}(\mathbb{C})$ with at most logarithmic growth as $|z|\to\infty$ and $|\nabla f(z)|\leqslant C(1+|z|)^{C}$, we denote the probability measure,
@@
We now proceed to apply the Ward identity to the functions $h_{1}$, $h_{2}$ and $h_{3}$. Let $g_{\infty}=\lim_{|z|\to\infty}g_{\rm out}(z)$. Using $\bar{\partial}g_{+}=\partial g_{-}=0$ on $S^{\rm c}$, by Green’s theorem we obtain:
$\int\frac{h_{1}(w)}{z-w}{\rm d}\mu_{V}=\frac{1}{\pi}\int_{\mathbb{C}}\frac{\bar{\partial}g_{+}}{z-w}=g_{+}(z)-g_{\infty},\quad\int\frac{h_{2}(w)}{z-w}{\rm d}\mu_{V}=\frac{1}{\pi}\int_{\mathbb{C}}\frac{\overline{\partial g_{-}}}{z-w}=\overline{g_{-}(z)}$
@@
Let $C,\kappa>0$ be fixed constants. Let $g=g_{1}+g_{2}$ where $g_{1}\in\mathscr{S}_{3,C,\kappa}$, $g_{2}\in\mathscr{C}^{3}(\mathbb{C})$ supported in $S_{\rm out}(1/2)$ with at most logarithmic growth as $|z|\to\infty$ and $\|\nabla g_{2}\|_{\infty}+\|\nabla^{2}g_{2}\|_{\infty}\leqslant C(1+|z|)^{C}$. Let $f$ be of class $\mathscr{C}^{5}(\mathbb{C})$ with at most logarithmic growth as $|z|\to\infty$ and $|\nabla f(z)|\leqslant C(1+|z|)^{C}$. Then, for any $D>0$,
@@
Note that if no cancellation were expected, with overwhelming probability, the number of indices $i$ with $g(z_{i})\asymp 1$ would be $\asymp\tau^{2}N$. For each such $i$, the sum over $j$’s for which $\frac{\overline{z_{i}}-\overline{z_{j}}}{z_{i}-z_{j}}e^{-\frac{|z_{i}-z_{j}|^{2}}{\delta^{2}}}\asymp 1$ would contribute about $N^{2\kappa}$. Consequently, one would expect the total sum to be on the order of $\tau^{2}N^{1+2\kappa}$. The lemma above shows, however, that angular cancellations occur and the apparent leading contribution is substantially reduced. Some angular cancellations were also instrumental in [LebSer2018, BauBouNikYau2019], in which isotropy was obtained differently from partition function asymptotics, for arbitrary inverse temperature. The above lemma is specific to the free fermion setting, and the proofs proceeds very differently, to cover cancellations up to the edge of the droplet, and for general biased measures.
@@
First step: reducing to the bulk. Let $\chi$ be a smooth bump function that is equal to $1$ on $S_{\rm in}(2N^{2\kappa}\delta)$, and 0 outside $S_{\rm in}(N^{2\kappa}\delta)$. We first consider the sum,

### v2

@@ 1 Introduction
The main Theorem 1.1 below generalizes the above moment formula to the exponential generating function of the electric potential, coupled with mesoscopic linear statistics, for general confining external potentials. A remarkable aspect of this general formula is the relevance of the harmonic measure on the boundary of the droplet, with the following consequence: the Gibbs measure associated to the electric potential converges to a Gaussian multiplicative chaos measure, which is universal up to a global Gaussian shift inherited from the capacity of the droplet (see Theorem 1.7).
@@
where the Barnes function $\rm G$} is defined in Subsection 1.5. Applications of general Fisher-Hartwig asymptotics for Toeplitz determinants (including the case of jump singularities) are multiple, we refer for example to [DeiItsKraII] for the relevance to phase transitions in the Ising model. The main contribution of our paper is the analogue of (1.2) – a formula due to Harold Widom in 1973 – for non-Hermitian random matrices, as explained below.
@@
where ${\rm d}M$ is the measure induced by the Lebesgue measure on all $N\times N$ complex matrices and $V:\mathbb{C}\to\mathbb{R}\cup\{+\infty\}$ be a confining potential. This induces the following probability measure on the eigenvalues of $M$:
@@
Before stating it, we introduce some notation and recall some facts from potential theory (see e.g. [GarMar2005, Chapters II-III]). For a function $g$ with domain including $\bar{S}$, let $g^{S}$ denote the unique bounded harmonic extension of $g$ to $S^{\rm c}$, with $g^{S}=g$ on $\bar{S}$. Note that for this harmonic extension, $g^{S}(\infty):=\lim_{|z|\to\infty}g^{S}(z)$ exists and
$|g^{S}(z)-g^{S}(\infty)|=\operatorname{O}\big(|z|^{-1}\big),\quad|\nabla^{n}g^{S}(z)|=\operatorname{O}\big(|z|^{-n-1}\big)\quad\textnormal{for all}\ n\geqslant 1\ \textnormal{as}\ |z|\to\infty.$
@@
Denote $\log^{\zeta}=\log|\zeta-\cdot\,|$. The Green’s function with a pole at $\infty$ on the exterior domain $S^{\rm c}$ is defined by (see e.g. [GarMar2005, Chapter II.2])
$G(z,\infty)=\log^{\zeta}z-(\log^{\zeta})^{S}(z),\quad\zeta\in S$
where the right hand side is independent of the choice of $\zeta\in S$ (see Appendix B) and we omitted the dependence on $S$ from the notation. The capacity of the set $\bar{S}$ is defined by (see e.g. [GarMar2005, Chapter III.1])
$\mathfrak{s}:=\log\operatorname{cap}(\bar{S})=(\log^{\zeta})^{S}(\infty)$ (1.4)
where the right hand side is again independent of the choice of $\zeta\in S$ (see Appendix B). Equivalently, $\log\operatorname{cap}(\bar{S})$ is the infimum of the energy integral
$I(\mu)=\iint\log\frac{1}{|z-w|}{\rm d}\mu(z){\rm d}\mu(w)$ (1.5)
over all probability measures $\mu$ supported on $\bar{S}$, [GarMar2005, Thm. III.4.1]. The infimum is uniquely attained by the harmonic measure seen from infinity which is supported on the boundary $\partial S$ and given by
${\rm d}\omega^{\infty}(z)=\frac{1}{2\pi}\frac{\partial G(z,\infty)}{\partial n}{\rm d}s=\frac{1}{2\pi}\mathcal{N}(\log^{\zeta})(z){\rm d}s,\quad z\in\partial S$ (1.6)
where ${\rm d}s$ denotes arc-length measure on $\partial S$, [GarMar2005, Thm. II.2.5]. See Appendix B for a self-contained proof of the independence of the right-hand side from the choice of $\zeta\in S$ and the fact that $\omega^{\infty}$ is a probability measure. Equivalently, $\omega^{\infty}$ is characterized probabilistically as follows. Let $B^{z}$ be planar Brownian motion started from $z\in\mathbb{C}$, and $\tau$ denote the hitting time of $\partial S$. Then, $\omega^{\infty}$ coincides with the weak limit of the law of $B^{z}_{\tau}$ as $|z|\to\infty$. Finally, for any $\zeta\in S$, the Poisson-Jensen formula applied to $(\log^{\zeta})^{S}$ yields
$\mathfrak{s}=(\log^{\zeta})^{S}(\infty)=\int_{\partial S}\log^{\zeta}(z)\,{\rm d}\omega^{\infty}(z),\quad\textnormal{for all}\ \zeta\in S.$ (1.7)
We also need the notations $L=\log\frac{\Delta V}{4}$, and refer to Subsection 1.5 for the definition of the functional space $\mathscr{S}_{n,C,\kappa}$, to state the following main theorem. It gives the asymptotics for the joint Laplace transform of the electrostatic potential, or equivalently the joint moments of the characteristic polynomial of random normal matrices.
@@
Let $C>1$, $m\in\mathbb{N}$, $0<\kappa\leqslant\frac{1}{100C}$ be fixed constants. Then, uniformly in $f\in\mathscr{S}_{5,C,\kappa}$, $\zeta_{1},\dots,\zeta_{m}\in\{\zeta\in S:\operatorname{dist}(\zeta,\partial S)>N^{-1/2+\kappa}\}$ satisfying $\min_{1\leqslant j<k\leqslant m}|\zeta_{j}-\zeta_{k}|>N^{-1/2+\kappa}$ and $\gamma_{1},\dots,\gamma_{m}\in[0,C]$ we have ($\rm G$} stands for the Barnes $\rm G$}-function, see Subsection 1.5)
$\mathbb{E}\Big[e^{\sum_{i=1}^{N}f(z_{i})}\prod_{j=1}^{m}\prod_{i=1}^{N}|z_{i}-\zeta_{j}|^{\gamma_{j}}\Big]=e^{N\int f{\rm d}\mu_{V}+\frac{1}{8\pi}\int_{\mathbb{C}}(|\nabla f^{S}|^{2}+\mathds{1}_{S}\Delta f+\Delta fL^{S}){\rm d}m}\prod_{j=1}^{m}e^{\frac{\gamma_{j}}{2}\big(\int_{\partial S}f{\rm d}\omega^{\infty}-f(\zeta_{j})\big)}$
$\times\prod_{j=1}^{m}e^{\gamma_{j}N\int\log^{\zeta_{j}}{\rm d}\mu_{V}}N^{\frac{\gamma_{j}^{2}}{8}}\frac{(2\pi)^{\frac{\gamma_{j}}{4}}}{{\rm G}\left(1+\frac{\gamma_{j}}{2}\right)}e^{\frac{\gamma_{j}}{4}(L(\zeta_{j})-L^{S}(\infty))+\frac{\gamma_{j}^{2}}{8}(L(\zeta_{j})+2\mathfrak{s})}$
$\times\prod_{1\leqslant j<k\leqslant m}|\zeta_{j}-\zeta_{k}|^{-\frac{\gamma_{j}\gamma_{k}}{2}}e^{\frac{\gamma_{j}\gamma_{k}}{2}\mathfrak{s}}\left(1+\operatorname{O}(N^{-\kappa/10})\right).$
The Ginibre ensemble is the paradigmatic non-Hermitian random matrix model; see [ByuFor] for its many facets. For this ensemble, our asymptotics take a simpler form. We have $\int_{\mathbb{D}}\log^{\zeta}\frac{{\rm d}m}{\pi}=\frac{|\zeta|^{2}-1}{2}$ and choosing $\zeta=0$ gives $\mathcal{N}(\log^{\zeta})=1$ and $\int_{\partial\mathbb{D}}\log^{\zeta}{\rm d}s=0$, so $\log\operatorname{cap}(\bar{\mathbb{D}})=0$. Hence the main theorem takes the following simplified expression, conjectured by Webb and Wong in [WebWong2019].
@@
To conclude this subsection, we note that moments of random characteristic polynomials of wide classes of random matrices have been a topic of major interest, see e.g. [AkeVer, FyoSup, BorStr, ForRai2009, BaiKea, Afa2019, SerSim2025] in the case of integer exponents by algebraic and supersymmetric methods, and [Kra2007, ItsKra, Cha, BerWebWon2017, WebWong2019, ChaFahWebWon, CFLW2021, ByuSeoYan, DeaMclMolSim2025] for fractional exponents by Riemann-Hilbert methods. Motivations for the study of random determinants in the non-Hermitian setting include for example the stability of large complex systems [BenFyoKho2021]. Theorems 1.1 and 1.2 introduce joint fractional moments for non-Hermitian ensembles, motivated by two-dimensional random geometry from free fermions, as explained below.
@@
In the 2d ambient space the Liouville quantum gravity measure also appears as stated in Theorem 1.7 below. The key ingredient is the precise asymptotics of the moments of the characteristic polynomial, Theorem 1.1 above. To state our result, note that from Corollary 1.6, the log-characteristic polynomial of a random normal matrix is $\frac{1}{2}$-log-correlated, so the corresponding critical value becomes $\gamma_{\rm c}=2\sqrt{2}$. The regime $\gamma<2$ is often referred to as the $L^{2}$-phase, while the full subcritical range $\gamma<2\sqrt{2}$ is called the $L^{1}$-phase. All arguments below are carried out in the $L^{1}$-phase.
@@
$\mathbb{E}\big[\log|\det(M-z)|\big]=N\int\log^{z}{\rm d}\mu_{V}+\frac{1}{4}+\frac{1}{4}(L(z)-L^{S}(\infty))+\operatorname{O}(N^{-\varepsilon}).$
@@
where the second equality is shown in Appendix B. Note that $\mathfrak{s}=\log\operatorname{cap}(\bar{S})$ is the smallest real number for which the kernel $T$ is positive semi-definite, since $\log\operatorname{cap}(\bar{S})$ is the infimum of the energy integral given in (1.5).
Let $\nu_{\gamma,(\mathfrak{s})}$ be the GMC measure associated to $X$, formally given by
@@
Our main theorems also allow to recover the leading order asymptotics for the electric field for general potential: For $K$ any compact subset of $S$ with non-empty interior and any $\varepsilon>0$,
@@
Steps (2) and (4) rely on the method of Ward identities, or loop equations, which is common in field theories. For Hermitian random matrices, the seminal work [Joh1998] relied on loop equations to reduce Laplace transform estimates to a shift of the density of states. In our non-Hermitian setting, Ward identities are also instrumental as in [AmeHedMak2015], but the control of the error due to potential anisotropy (i.e. local directional contribution from the interactions) and the boundary of the droplet, is delicate for measures biased by determinant powers, i.e. in the presence of root singularities. Regarding the proof of local isotropy – a problem naturally absent in the settings of [BouFal2025, Keles2025] – Lemma 2.14 first treats nonsingular potentials, by a moments method specific to determinantal point processes but flexible enough to cover mesoscopic observables up to the boundary of the droplet; this estimate is then transferred to singular setting, as it holds with overwhelming probability (here $1-e^{-(\log N)^{D}}$ for any $D>0$), while in [AmeHedMak2015] angular cancellations hold in expectation for the nonsingular potentials. About the effect of the boundary, one novelty of our analysis is a careful control of the norms of the harmonic extensions of functions on mesoscopic scales, and an analysis of the effect of bulk singularities on the boundary field. This analysis is significant due to the harmonic measure on the boundary of the droplet, which generates an additional Gaussian shift when compared to the free field with free boundary conditions, harmonically extended outside of the droplet.
@@ Acknowledgements.
We wish to thank Fredrik Viklund for enlightening comments about harmonic measures. P. B. and A. K. were supported by the NSF standard grant DMS-2348202. G. D. gratefully acknowledges support from the Fondation de l’École Polytechnique, as well as from the Agence Nationale de la Recherche (ANR-25-CE40-5672 and ANR-25-CE40-1380). The research of L. H. was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) in the framework to the collaborative research center ”Multiscale Simulation Methods for Soft-Matter Systems” (TRR 146) under Project No. 233630050 and Project-No. 443891315 within SPP 2265.
@@ 2 Preliminaries on nonsingular potentials
For any given function $f$ of class $\mathscr{C}^{5}(\mathbb{C})$ with at most logarithmic growth as $|z|\to\infty$ and $|\nabla f(z)|\leqslant C(1+|z|)^{C}$ (for a possibly $N$-dependent $C$), we denote the probability measure
@@
We now proceed to apply the Ward identity to the functions $h_{1}$, $h_{2}$ and $h_{3}$. Recall that we denote $\lim_{|z|\to\infty}g_{\rm out}(z)$ by $g^{S}(\infty)$. Using $\bar{\partial}g_{+}=\partial g_{-}=0$ on $S^{\rm c}$, by Green’s theorem we obtain:
$\int\frac{h_{1}(w)}{z-w}{\rm d}\mu_{V}=\frac{1}{\pi}\int_{\mathbb{C}}\frac{\bar{\partial}g_{+}}{z-w}=g_{+}(z)-g^{S}(\infty),\quad\int\frac{h_{2}(w)}{z-w}{\rm d}\mu_{V}=\frac{1}{\pi}\int_{\mathbb{C}}\frac{\overline{\partial g_{-}}}{z-w}=\overline{g_{-}(z)}$
@@
Let $C,\kappa>0$ be fixed constants. Let $g=g_{1}+g_{2}$ where $g_{1}\in\mathscr{S}_{3,C,\kappa}$, $g_{2}\in\mathscr{C}^{3}(\mathbb{C})$ supported in $S_{\rm out}(1/2)$ with at most logarithmic growth as $|z|\to\infty$ and $\|\nabla g_{2}\|_{\infty}=\operatorname{O}(N^{\log\log N})$. Let $f$ be of class $\mathscr{C}^{5}(\mathbb{C})$ with at most logarithmic growth as $|z|\to\infty$ and $|\nabla f(z)|\leqslant c(1+|z|)^{c}$ for an $N$-dependent constant $c$. Then, for any $D>0$,
@@
Note that if no cancellation were expected, with overwhelming probability, the number of indices $i$ with $g(z_{i})\asymp 1$ would be $\asymp\tau^{2}N$. For each such $i$, the sum over $j$’s for which $\frac{\overline{z_{i}}-\overline{z_{j}}}{z_{i}-z_{j}}e^{-\frac{|z_{i}-z_{j}|^{2}}{\delta^{2}}}\asymp 1$ would contribute about $N^{2\kappa}$. Consequently, one would expect the total sum to be on the order of $\tau^{2}N^{1+2\kappa}$. The lemma above shows, however, that angular cancellations occur and the apparent leading contribution is substantially reduced. Some angular cancellations were also instrumental in [LebSer2018, BauBouNikYau2019], in which isotropy was obtained differently from partition function asymptotics, for arbitrary inverse temperature. The above lemma is specific to the determinantal setting, and the proofs proceeds very differently, to cover cancellations up to the edge of the droplet, and for general biased measures.
@@
First step: reducing to the bulk. Let $\chi$ be a smooth bump function that is equal to $1$ on $S_{\rm in}(2N^{2\kappa}\delta)$, and 0 outside $S_{\rm in}(N^{2\kappa}\delta)$. We first consider the sum